# Towards Better Open-Ended Text Generation: A Multicriteria Evaluation Framework

### v2

@@ 4 Benchmarking utilizing Order Structure
Metric information considered Order only Order only Order and metric value
Methodology basis Pairwise comparison Partial orders Mean values
@@
Output Returns a worth parameter and a total order Returns a partial order Returns mean values and a total order
@@
Results (based on WikiText-103) Mistral 3 CS ((’0.4’, ’10’)) has the highest worth parameter, while GPT2-XL CS ((’1.0’, ’20’)) has the lowest. The top five models in the Extended Bradley-Terry Model are incomparable, despite the model suggesting a total order. Falcon 2 CS ((’0.8’, ’1’)) has the highest mean and Mistral 3 CS ((’0.2’, ’1’)) the lowest.
@@
Table 2: Comparison of the three methods: extended Bradley-Terry, ufg-depth and Q*Text, see also Figure 1.
To benchmark decoding methods with respect to multiple criteria (see Section 2) aiming for a ranking of methods (see practitioners example in Section 1), we adopt very recent developments in the theory of multicriteria and multi-task benchmarking Cruz et al. (2024); Zhang and Hardt (2024b); Kohli et al. (2024); Jansen et al. (2024, 2023b, 2023a); Rodemann and Blocher (2024); Blocher et al. (2024), some of them grounded in decision theory (social choice theory), some in the theory of data depth. The starting point of all of them is a self-evident problem: When comparing algorithms with respect to multiple criteria (e.g., metrics or tasks) on multiple test problems (e.g., datasets), one algorithm can be better than another one with respect to one criterion while being worse with respect to another criterion. Aggregating such results into a single ranking is typically either done via the Pareto-front (all algorithms that are undominated with respect to all criteria) or by weighted sums of the criteria. While the former is hardly informative for large-scale benchmarking, the latter arbitrarily depends on the choice of weights.
In this section, we propose benchmarking of decoding strategies along the Bradley-Terry model Bradley and Terry (1952b) and the union-free-generic (ufg) depth Blocher et al. (2024); Blocher and Schollmeyer (2024). Both approaches deliver rankings of decoding strategies rather than metric values for each strategy. We motivate this from a very practical perspective, see Section 1: In many cases, the metric information can be considered redundant. That is, ranking the decoding strategies is the overall aim of benchmarking, not assigning metric values. After all, a decoding method can either be deployed or not by practitioners, rendering the metric information not of direct practical interest.
4.1 Extended Bradley-Terry Model
The extended Bradley-Terry model is based on pairwise comparisons Bradley and Terry (1952a); Davidson (1970). To obtain this pairwise comparison of the methods, we compare the methods pairwise on each single generation. For simplicity, let us look at WikiText-103 with 1314 generations. For each generation $G_{i,j}$ (from model $i$ with decoding strategy $j$, see Section 1), we can evaluate the methods based on coherence, perplexity, and diversity. Now we consider each pair of methods and each generation, e.g. Mistral 3 CS ((’0.2’, ’1’)) and Mistral 3 CS ((’0.8’, ’1’)) with generation id 1. Then we state that method 1 dominates method 2 if and only if there exists one metric that strictly prefers method 1 over method 2 and the other two metrics do not disagree. Based on generation id 1, we obtain that Mistral 3 CS ((’0.2’, ’1’)) is preferred over Mistral 3 CS ((’0.8’, ’1’)) since coherence states that method Mistral 3 CS ((’0.2’, ’1’)) is preferred over method Mistral 3 CS ((’0.8’, ’1’)), and the other two metrics, here diversity and perplexity, do not contradict this statement. In this way, for WikiText-103 we get 1314 comparison statements for each pair of two methods. Note that when two methods have exactly the same metric values, they are considered indifferent. This differs from incomparability; for details, see Rodemann and Blocher (2024). However, with slight abuse, we do not distinguish these two cases here.
The extended Bradley-Terry model returns so-called "worth" parameters for each method, which indicate the probability that one model is preferred over another in a pairwise comparison. Thus, the method with the highest worth dominates every other method. Since it is unlikely that two worth parameters will have exactly the same value, the extended Bradley-Terry model results in a total order. Thus, it sets any two methods to be in relation to each other and therefore it does not contain any information on the incomparability and indifference of two methods. In light of our previous considerations, enforcing a strict total order of methods may be a too strong assumption that does not reflect the underlying complexity of method evaluation.
4.2 Union-Free Generic Depth
To account for exactly this complexity, we build on Rodemann and Blocher (2024); Blocher et al. (2024) and deliberately allow for incomparability. Mathematically, this means that we aim for partial rather than total orders. Let us look again at a single generation $G_{i,j}$. For the extended Bradley-Terry model, we only considered the pairwise comparisons. However, all the pairwise comparisons resulting from a generation $G_{i,j}$ together define a partial order that describes the preference structure of the methods based on that generation $G_{i,j}$. This gives us 1314 partial orders for the WikiText-103 data. For example, in the case where we compare four methods, the two partial orders in Figure 3 correspond to two observation.
The union-free generic depth computes a measure for each partial order that indicates how central/typical it is compared to the observed partial orders. Thus, it states how central/typical a preference structure of all considered methods in WikiText-103 is. This allows us to identify the most central ranking, i.e. the ranking that is most supported by the observed data.
5 Benchmarking by Collapsing into a Single Metric
Our multicriteria analysis in Section 4 offered practical insights into rankings among decoding methods on different tasks. While in many applications these types of rankings might be sufficient, other scenarios require a single, unified metric. Consider, for instance, again the example in Section 1 of a researcher aiming to gain general insights into how much better one method is compared to another one. Moreover, aggregating all criteria into a single metric simplifies the evaluation process and aids in making more definitive comparisons between different decoding methods. In this section, we introduce a method for collapsing the multicriteria evaluation into a singular metric, Q*Text, providing a streamlined approach for comprehensive assessment.
As previously discussed, metrics such as coherence and diversity, when optimized excessively, can lead to unintended outcomes. High coherence scores often result in repetitive or degenerate text, as the model overemphasizes continuity at the expense of introducing new information. Conversely, prioritizing high diversity can produce erratic sequences that lose relevance to the initial prompt, undermining the overall quality of the generated text. Q*Text consists of the harmonic mean of normalized and penalized coherence, diversity and generation perplexity. Its construction involves three key steps: normalization to a common scale of 0 to 100, where 0 represents the worst performance and 100 represents the best, penalization of extreme values, and aggregation using the harmonic mean. A technical description is detailed in Section A.2.
6 Results
In this section, we present our main results. First, we look at the extended Bradley-Terry model and the union-free generic depth approach, both of which are based on pairwise comparison. A first impression can be seen in Appendix A.3. We observe that out of a total of 124962 pairwise comparisons among methods, only a very small percentage shows a clear dominance structure. That is, for the majority of the method comparisons, the metrics contradict each other (or imply indifference), e.g., a method is better with respect to coherence but worse with respect to diversity. For a small minority of comparisons, however, methods multilaterally outperform other methods, e.g., method 1 outperforms method 2 for at least 90% of the generations with respect to all metrics.
Extended Bradley-Terry Model
@@ 6 Results
The extended Bradley-Terry model returns so-called "worth" parameters which indicate the probability that this decoding method is preferred over the other in a pairwise comparison. When all datasets are considered at once, the method that dominates all other methods according to the extended Bradley-Terry model is Mistral 3 CS ((’0.6’, ’15’)). The second best method is Mistral 3 CS ((’0.4’, ’5’)). The worst method is GPT2-XL CS ((’1.0’, ’20’)). A snippet of the results, including the case when restricting the analysis to only one dataset, can be found in Table 3.
Method Estimated worth parameter
@@
Table 3: Estimated worth parameter of the extended Bradley-Terry model based on WikiText-103 dataset and the metric coherence, diversity and perplexity.
@@
Note that the total order provided by the extended Bradley-Terry model respects the pairwise dominance structures discussed in Appendix A.3. As noted above, the extended Bradley-Terry model leads to a total order and neglects information about incomparabilities. However, the dominance structure provided by the partial orders given by each generation, see Appendix A.3, already suggests that enforcing a total order (e.g., not allowing incomparability of two methods) may be too strong an assumption. Moreover, the extended Bradley-Terry model assumes that the comparison of method 1 with method 2 is independent of the comparison between method 3 and method 4, but this seems to be questionable for a fixed generation, see Blocher et al. (2024).
Union-Free Generic Depth
Therefore, in the next step, we consider the union-free generic depth approach, which allows for two methods to be incomparable. Furthermore, the ufg-depth considers the entire set of pairwise comparisons for a generation as one observation and does not assume an independence structure between them. Due to the high computational complexity, see Appendix A.7, we restrict our analysis to the WikiText-103 dataset and compare only the four methods that appear to be the best according to the extended Bradley-Terry model, see Appendix A.4: Mistral 3 CS ((’0.6’, ’15’)), Mistral 3 CS ((’0.4’, ’3’)), Mistral 3 CS ((’0.8’, ’3’)) and Mistral 3 CS ((’0.4’, ’20’)).
The highest ufg-depth with a value of 0.977 (thus the one that has the structure most supported by the observation), is the one that shows no dominance structure among the four methods, i.e. the one that concludes that all methods are incomparable to each other, see Figure 3 (top). Roughly speaking, our method reveals that the four decoding methods considered here are incomparable. More formally put, we identify a trivial ranking with no dominance structure as the “central” (in the sense of being the “median”) of the dataset comprising the benchmarking results. This means that such a ranking has most support by the benchmarking results. Our method further finds an “outlier”, i.e., a ranking of methods that has least support by the benchmarking results. In the example at hand, this outlier is a partial ranking that ranks Mistral 3 CS ((’0.4’, ’3’)) higher than Mistral 3 CS ((’0.4’, ’20’)), see Figure 3 (bottom). This means that given the benchmarking results, such a ranking of methods is “least central” or “atypical” and therefore based on the benchmarking results the least supporter structure.
@@
Figure 3: Partial orders with the highest (top) and lowest (bottom) ufg-depths based on the four methods and WikiText-103.
Q*Text
With regards to Q*Text, we observe that balanced deterministic strategies such as contrastive search with moderate penalties ($\alpha$ values of 0.4 or 0.6) and moderate values of $k$ (5, 10, or 15) are preferred, as illustrated in Tables 18 and 16. Combinations that counterbalance each other also yield favorable results – for instance, low $\alpha$ values (0.2) with high $k$ values (20 or 50), or high $\alpha$ values (0.8 or 1.0) with moderate $k$ values (3 or 5). Both Q*Text and the extended Bradley-Terry model indicate a general preference for larger architectures; however, smaller models like GPT2-XL can outperform more modern architectures when paired with balanced decoding strategies, see Section 12.
Regarding stochastic methods, Q*Text shows a preference for strategies that enhance diversity. This includes sampling with temperatures above 0.7, top-$k$ sampling with $k$ values exceeding 10, and nucleus (top-$p$) sampling with $p$ values above 0.8. Conversely, beam search is generally disfavored due to its extremely low diversity.
To evaluate the level of agreement between the total ordering from the extended Bradley-Terry and that provided by Q*Text, we analyze and visualize the decoding methods with the highest agreement and those with the largest rank deviations (see Appendix A.6, Figures 4 and 5). The results suggest discrepancies, particularly for less diverse and less coherent generations. Nevertheless, there is good agreement when ranking decoding methods with moderate hyperparameters, as previously discussed.
Despite operating on different underlying mechanisms, the preferences from the extended Bradley-Terry model exhibit noteworthy similarities with those from Q*Text concerning the aforementioned balanced strategies. One notable difference is that the extended Bradley-Terry model does not account for significant drops in diversity, as seen with beam search, which reduces the penalization of this decoding strategy. On the other hand, decoding methods such as GPT2-XL with contrastive search parameters ($\alpha$ of 1.0 and $k$ of 20) are strongly penalized by both approaches due to their incoherence and low generation confidence, see Tables 13, 15 and 19.
Finally, we naturally ask how these alternative methods, evaluated within a multicriteria framework, align with human preferences. To address this, we considered the rankings provided by human evaluators in Section 5 of Garces-Arias et al. (2024a), which present similar findings. Specifically, human evaluators favored balanced contrastive search strategies, e.g., (0.4, 10), (0.4, 20), (0.4, 5), (0.6, 10), and (0.6, 5), top-$k$ sampling with $k=10$ or 50, and nucleus (top-$p$) sampling with $p=0.9$. In contrast, unbalanced contrastive search approaches, such as (1.0, 50) and (0.8, 50), were consistently evaluated as low quality due to their low coherence. Beam search strategies with moderate beam widths, e.g. less than or equal to 20, were also often judged as low quality due to their low diversity. See Tables 10, 11 and 18 and the work by Garces-Arias et al. (2024a) for more details. Furthermore, Table 20 presents examples of generations from various decoding methods, highlighting those that received positive and negative evaluations from Bradley-Terry and Q*Text.

### v3

@@ 4 Scenario 1: Ranking Methods
Considered Information Order only Order only Order and metric value
Methodology Pairwise comparison Partial orders Mean values
@@
Output Worth Parameter & Total Order Partial Order Mean Values & Total Order
@@
Results (WikiText-103) Mistral 3 CS ((’0.4’, ’10’)) has the highest worth parameter, while GPT2-XL CS ((’1.0’, ’20’)) has the lowest The top five models in the Extended Bradley-Terry Model are incomparable, despite the suggested total order Falcon 2 CS ((’0.8’, ’1’)) has the highest mean and Mistral 3 CS ((’0.2’, ’1’)) the lowest
@@
Table 2: Comparison of the extended Bradley-Terry Model, the ufg-depth and Q*Text (cf. Figure 1).
4.1 Extended Bradley-Terry Model: Theory
The extended Bradley-Terry model is based on pairwise comparisons Bradley and Terry (1952a); Davidson (1970). It offers a flexible way to rank items while respecting both clear dominance structures and non-dominances (i.e., ties). Each item $i$, in our situation, decoding method $i$, is assigned a worth parameter $\pi_{i}$. These worth parameters represent the relative performance/strength of a decoding method in comparison to another decoding method, with all worth parameters summing up to one. The probability that decoding method $i$ is preferred over decoding method $j$ is $P(i>j)=\pi_{i}/(\pi_{i}+\pi_{j}+\nu\sqrt{\pi_{i}\pi_{j}})$. Here, $\nu$ is a discrimination parameter that reflects the likelihood of a tie, i.e., no preference between the two decoding methods. Based on the estimations, it is possible to conclude that decoding methods with high worth parameters dominate others.
Sinclair (1982) reformulated the extended Bradley-Terry model as a generalized linear model (GLM) with a Poisson distribution and log link: Let $m_{i>j}$ be the count of times decoding method $i$ outperforms decoding method $j$ and $m_{i\sim j}$ be the number of ties. Then the GLM is given by $\log(m_{i>j})=\mu_{ij}+\frac{1}{2}\log(\pi_{i})-\frac{1}{2}\log(\pi_{j})$ and $\log(m_{i\sim j})=\mu_{ij}+\log(\nu)$ with parameters $\mu_{ij}=\ln{m}-\ln{\left(\sqrt{\pi_{i}/\pi_{j}}+\sqrt{\pi_{j}/\pi_{i}}\right)}$ and $m$ the total number of pairwise comparisons.
Since it is unlikely that two worth parameters have exactly the same value, the extended Bradley-Terry model yields a total order representing the performance of the decoding methods across all prompts.
4.2 Extended Bradley-Terry Model: Experimental Results
@@
The extended Bradley-Terry model returns so-called "worth" parameters, which indicate the probability that this decoding method is preferred over the other in a pairwise comparison. When all datasets are considered at once, the method that dominates all other methods according to the extended Bradley-Terry model is Mistral 3 CS ((’0.6’, ’15’)). The second-best method is Mistral 3 CS ((’0.4’, ’5’)), while the worst method is GPT2-XL CS ((’1.0’, ’20’)). An excerpt of the results, including the case when restricting the analysis to only one dataset, is presented in Table 3.
Decoding Method Estimated worth parameter
@@
Table 3: Estimated worth parameter of the extended Bradley-Terry model based on WikiText-103 dataset, and the metrics coherence, diversity and perplexity.
@@
Note that the total order provided by the extended Bradley-Terry model respects the pairwise dominance structures discussed in Appendix C. As noted above, the extended Bradley-Terry model leads (in almost all cases) to a total order. Hence, it neglects information about incomparabilities. However, the dominance structure provided by the partial orders given by each generation, see Appendix C, already suggests that enforcing a total order (e.g., not allowing incomparability of two decoding methods) may be too strong an assumption. Additionally, the extended Bradley-Terry model relies on further independence assumptions that may not be appropriate for benchmarking purposes (Blocher et al., 2024).
4.3 Union-Free Generic Depth: Theory
The union-free generic (ufg) depth (Rodemann and Blocher, 2024; Blocher et al., 2024) directly addresses these concerns by incorporating incomparability information in the estimation itself and avoids any additional independence assumptions. Mathematically, this means that we aim for partial rather than total orders. Let us look again at a single prompt and the procedure discussed directly before Section 4.1. For the extended Bradley-Terry model, we only considered the pairwise comparisons. However, all the pairwise comparisons resulting from one single prompt define a partial order that describes the performance of the decoding methods based on that single prompt. This yields 1,314 partial orders for the WikiText-103 data. For example, in the case where we compare four decoding methods, the two partial orders in Figure 2 correspond to two observations.
@@ 4.3 Union-Free Generic Depth: Theory
Figure 2: Partial orders with the highest (top) and lowest (bottom) ufg-depths based on Wikitext-103 and the four decoding methods presented in Table 3
The ufg-depth analysis provides a measure for each partial order that indicates how central/typical or outlying/atypical it is. Since each partial order represents the performance of the decoding method, the ufg-depth provides insights into typical and atypical performance structures of the decoding methods. This allows us to identify the most central ranking, i.e., the ranking that is most supported by the observed data. To achieve this, the ufg-depth generalizes the well-known simplicial depth from $\mathbb{R}^{d}$ (which measures centrality by the probability that a point $x$ lies in a randomly drawn $d+1$ simplex Liu (1990)) to partial orders. This is, Blocher et al. (2024) generalize the meaning of "lying in" and "$d+1$ simplex" for $\mathbb{R^{d}}$, which can be defined by the convex closure operator and the convex sets, to partial orders. Let $\mathcal{P}$ be the set of all partial orders given by the items/decoding methods $m_{1},\ldots,m_{k}$. To transfer the idea of "lying in", Blocher et al. (2024) considered the closure operator $\gamma:2^{\mathcal{P}}\rightarrow 2^{\mathcal{P}},P\mapsto\{p\in\mathcal{P}\mid\cap_{\tilde{p}\in P}\>\tilde{p}\subseteq p\subseteq\cup_{\tilde{p}\in P}\}$. Blocher et al. (2024) showed that $d+1$ simplices in $\mathbb{R}^{d}$ are those convex sets that are non-trivial, minimal, and not decomposable with respect to the convex closure operator. This is equivalent to consider those sets of partial orders $P=\{p_{1},\ldots,p_{k}\}\in\mathcal{S}\subseteq 2^{\mathcal{P}}$ that satisfy (I) $P\subsetneq\gamma(P)$ and (II) there exists no family $\left(B_{i}\right)$, with $i\in I$ index, such that $B_{i}\subseteq P$ and $\gamma(P)=\cup_{I}\gamma\left(B_{i}\right)$ (i.e. $P$ cannot be decomposed). The ufg-depth of a partial order $p$ is then the probability that $p$ lies in a randomly drawn $P\in\mathcal{S}$, weighted by the cardinality $P$, see Appendix B for details. For the empirical counterpart, we use the empirical probability measure.
4.4 Union-Free Generic Depth: Experimental Results
Therefore, in the next step, we consider the union-free generic depth approach, which allows for two methods to be incomparable. Furthermore, the ufg-depth considers the entire set of pairwise comparisons for a generation as one observation and does not assume an independence structure between them. Due to the high computational complexity, we restrict our analysis to the WikiText-103 dataset and compare only the four methods that appear to be the best according to the extended Bradley-Terry model, see Appendix D: Mistral 3 CS ((’0.6’, ’15’)), Mistral 3 CS ((’0.4’, ’3’)), Mistral 3 CS ((’0.8’, ’3’)) and Mistral 3 CS ((’0.4’, ’20’)).
The highest ufg-depth with a value of 0.977 (thus the one that has the structure most supported by the observation), is the one that shows no dominance structure among the four methods, i.e. the one that concludes that all methods are incomparable to each other, see Figure 2 (top). Roughly speaking, our method reveals that the four decoding methods considered here are incomparable. More formally put, we identify a trivial ranking with no dominance structure as the “central” (in the sense of being the “median”) of the dataset comprising the benchmarking results. This means that such a ranking has most support by the benchmarking results. Our method further finds an “outlier”, i.e., a ranking of methods that has least support by the benchmarking results. In the example at hand, this outlier is a partial ranking that ranks Mistral 3 CS ((’0.4’, ’3’)) higher than Mistral 3 CS ((’0.4’, ’20’)), see Figure 2 (bottom). This means that, given the benchmarking results, such a ranking of methods is “least central” or “atypical” and therefore based on the benchmarking results with the least supportive structure.
5 Scenario 2: Cardinal Assessment
While multicriteria analysis provides ordinal rankings among decoding methods, many applications require a single unified metric for benchmarking and optimization.
Use Case
We compute Q*Text scores for over 1.8M text continuations, as described in Table 1, and analyze their performance on a model level, decoding strategy level, and hyperparameter configurations level.
5.1 Q*Text: Theory
We propose Q*Text, a text quality metric that integrates coherence, diversity, and perplexity using weighted combinations with Gaussian penalty functions to handle extreme values.
Metric Formulation
Q*Text is defined as:
[…]
where $M_{i}$ are normalized metrics, $w_{i}$ are weights, and $P_{i}(x)=\exp(-\alpha_{i}(x-\mu_{i})^{2})$ are Gaussian penalties that discourage extreme values. Parameters $\mu_{i}$ represent optimal targets while $\alpha_{i}$ controls penalty strength.
Normalization
We apply inverse normalization to perplexity (lower is better): $M_{1}=\frac{p_{\max}-p_{i}}{p_{\max}-p_{\min}}$, and standard min-max normalization to coherence and diversity (higher is better): $M_{j}=\frac{m_{j}-m_{\min}}{m_{\max}-m_{\min}}$ for $j\in\{2,3\}$.
Parameter Optimization
The nine parameters $\theta=\{w_{i},\mu_{i},\alpha_{i}\}_{i=1}^{3}$ are optimized via:
$\theta^{*}=\text{argmax}_{\theta}\rho_{s}(\text{Q*Text}(\theta),H)$ (2)
where $\rho_{s}$ is Spearman correlation and $H$ are publicly available human ratings (Garces-Arias et al., 2025). The pseudo-code for the hyperparameter tuning of Q*Text, as well as an interpretation of the resulting values, are presented in Appendix G, Table 20, and Table 21. Finally, a visualization of the achieved $\rho_{s}$, highlighting alignments on a decoding strategy level, is illustrated in Appendix G, Figure 5.
5.2 Q*Text: Experimental Results
When analyzing the results we observe the following: For deterministic decoding methods, Q*Text favors balanced hyperparameter choices, particularly CS with moderate penalties ($\alpha$ values of 0.4 or 0.6) and moderate $k$ values (5, 10, or 15), as shown in Tables 16 and 18. Counterbalancing combinations also perform well, such as low $\alpha$ values (0.2) with high $k$ values (20 or 50), or high $\alpha$ values (0.8 or 1.0) with moderate $k$ values (3 or 5). Beam Search (BS) is generally disfavored due to extremely low diversity, indicating Q*Text’s capability to penalize degenerate text. For stochastic methods, Q*Text prefers diversity-enhancing strategies: temperature sampling with $\tau>0.7$, top-$k$ sampling with $k>10$, and nucleus sampling with $p>0.8$.
To illustrate specific results, we sample eight machine-generated continuations of a Wikitext prompt and include the original human text continuation. The text generations are produced by models of different sizes and decoding strategies with varying hyperparameter configurations. The results are presented in Table 4 and reveal a clear pattern: moderate decoding parameters produce reasonable continuations with scores ranging from 68 to 87, while extreme parameter settings lead to either repetitive or erratic text.
When the degeneration penalty reaches 1.0 with high top-$k$ values, models like GPT2-XL and Qwen 2 generate completely incoherent text with scores near zero. Similarly, Llama 3’s beam search produces repetitive, temporally inconsistent content. The human reference completion achieves a score of 87.33, establishing a quality benchmark. Interestingly, GPT2-XL with moderate parameters (degeneration penalty of 0.6, top-$k$ of 5) achieves 86.69, nearly matching human performance despite its relatively small size. These findings align with human evaluator preferences from Garces-Arias et al. (2025), who favored balanced CS strategies like ((’0.4’, ’10’)), ((’0.4’, ’20’)), ((’0.6’, ’10’)), top-$k$ sampling with $k=10$ or 50, and nucleus sampling with $p=0.9$. Unbalanced approaches like CS ((’1.0’, ’50’)) and BS with moderate beam widths were rated as low quality (Tables 11, 18).
Decoding Method Prompt: "Headlam served as Officer Commanding North @-@ Western Area in 1946, and as Director of Training from 1947 to 1950. In 1950 - 51," Q*Text
Deepseek CS (0.6, 15) "he was Chief of Staff (C.O.S.) of the 1st Parachute Brigade, and in 1951 – 52, he was G.O.C. of the 1st Parachute Brigade. He was promoted to Brigadier… (text continues)" 80.16
GPT2-XL CS (1.0, 20) "Headless Horsemurdertrud (his nom de veal et re, a translators and transcriph.coastavarism in Latin -Ceux du Ponto-Monahawak at Keremej, New Heythorhampsthembryghtes-sur Meriont Bay of Sulaunkhuyut/Arap Moheluwempeh is one hundred nautilus middelveth the… (text continues)" 0.02
Qwen 2 CS (1.0, 20) " Headelm oversa Operation Agulhene535af69, to reoccupyo portug 1 st colonial outst a ncesin AFRicacoun trie sto preva l sepa rabli st mov me s . In th es on he wro 4he couthoea2 rsh ipp ublicatin hitherto re t aine o secret… (text continues)" 0.02
Llama 3 beam (3) "he served as Commanding Officer of the 1st Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1951 2013 52, he served as Commanding Officer of the 2nd Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1952 2013 53, he served as Commanding Officer of the 3rd Battalion, The Queen’s Royal Regiment (West Yorkshire Regiment). In 1953 2013 54 , he served as Commanding Officer of the 4th Battalion, … (text continues)" 0.02
Table 4: Case Study: Comparison of multiple decoding methods for a prompt from the Wikitext corpus. The first five rows show examples generated by high-ranked methods, while the next three rows display those from low-ranked methods. Human-generated reference text is included for comparison. Degenerate text is highlighted in purple while erratic content is highlighted in brown.
6 Discussion
First, we examine the extended Bradley-Terry model and the union-free generic depth approach, both of which are based on pairwise comparisons. A first impression can be seen in Appendix C. We observe that out of a total of 124,962 pairwise comparisons among methods, only a very small percentage shows a clear dominance structure. That is, for the majority of the method comparisons, the metrics contradict each other (or imply indifference), e.g., a method is better with respect to coherence but worse with respect to diversity. For a small minority of comparisons, however, methods multilaterally outperform other methods, e.g., method 1 outperforms method 2 for at least 90% of the generations with respect to all metrics.
Moving on to Q*Text results, we observe that it shares a preference for larger architectures with the extended Bradley-Terry model, though smaller models like GPT2-XL can outperform modern architectures with balanced decoding strategies (Table 12).
Agreement analysis between the extended Bradley-Terry model and Q*Text (Appendix F, Figures 3 and 4) highlights discrepancies for less diverse and coherent generations, but good agreement for methods with moderate hyperparameters. The extended Bradley-Terry model does not penalize diversity drops as severely as Q*Text, while both approaches strongly penalize incoherent, low-confidence methods like GPT2-XL with CS ($\alpha=1.0$, $k=20$), see Tables 13, 15 and 19.
We now examine the advantages and disadvantages of the three proposed benchmarking methods within our established framework. As highlighted in Section 1, benchmarking serves different purposes: Scenario 1 requires only an ordering of decoding methods, while Scenario 2 additionally demands a cardinal assessment of quality. While Scenario 2 naturally encompasses Scenario 1, the ordering focus in Scenario 1 enables the utilization of partial ranking theory, leading to fundamentally different procedures than those based on mean transformations and incorporating concepts such as method incomparability.
Both Scenario 1 methods build upon a data transformation, where metric scores are translated into ordinal values. The extended Bradley-Terry Model offers computational efficiency with $O(n^{2}m)$ complexity, making it scalable to large numbers of methods and generations. It provides interpretable worth parameters representing estimated preference probabilities and addresses incomparabilities and ties in observed data. However, this approach forces a total order in results, potentially oversimplifying complex dominance structures where methods may genuinely be incomparable. The model assumes independence between pairwise comparisons, which is questionable when comparing methods on fixed datasets, and relies strictly on dominance agreements across all evaluated metrics.
The Union-Free Generic Depth method preserves incomparabilities through partial orderings, providing more realistic representations of method relationships while offering insights into entire performance distribution structures. Unlike the extended Bradley-Terry approach, it does not assume independence between pairwise comparisons, making it more suitable for fixed-dataset evaluations. Nevertheless, this method suffers from computational intensity with worst-case complexity $O(2^{m})$, limiting applicability to smaller methods and dataset subsets. The approach is more complex to interpret than traditional rankings and, like the extended Bradley-Terry method, may be overly conservative in establishing dominance relationships.
Q*Text provides cardinal assessment with meaningful score differences, enabling quantification of performance gaps. It incorporates penalization of extreme values to prevent degenerate solutions such as repetitive or erratic text, automatically balances multiple criteria through mean aggregation, and remains computationally efficient and straightforward to implement. However, the method relies on normalization bounds and penalization parameters that may not generalize across different contexts. By collapsing multiple metrics into a single score, it may obscure important trade-offs between individual metrics and prove less interpretable than separate metric examination, potentially masking insights about specific strengths and weaknesses.